From GlueXWiki
Jump to: navigation, search

March 20, 2014 FDC meeting


  1. Final installation and testing [1], [2]
    1. Testing with scope (Lubomir)
    2. Testing with DAQ E-log (Beni)
  2. Engineering (Bill, Dave)
    1. Cathode scan [3]
    2. Fixing tools [4]
  3. Electronics (Chris, Nick)
  4. Other


Participants: Eugene, Fernando, Bill, Chris, Nick, Simon, Beni, and Lubomir.

Final installation and testing

- FDC was inserted last Thursday. This time we used a cable support attached to the side rails with bearings. After insertion, connecting of the cooling system and turning on the LV we found one LV cable was disconnected in the third package. We borrowed a fancy HD videoscope from SRF and were able to look from the downstream side: probably one wire got caught on something and one side of the LV connector was disconnected. Bill and Dave prepared different tools (pictures attached) and on Monday we tried to insert the connector back but it didn't work. In the meantime we found 7 wire channels disconnected: later it turned out the signal cable was pulled and the connector was partially out.

- On Tuesday we took the FDC back on the platform. Casey fixed the LV connector; he had to fix also two pins on the daughter card connector that were bent. The wire channels and several cathode channels were also fixed. For 30 min. the TDCs were tested with DAQ and then FDC was inserted back (all done before noon). During the operation we kept the cooling connected and were monitoring the LV currents.

- After the first insertion Anatoly tested all the cathode channels with scope: about 50 bad channels to be compared to 18 before the insertion. After the re-insertion on Tuesday till now, Anatoly has tested half of the crates (we assumed Beni will be able to test the rest with DAQ) and there were no surprises. Anatoly will continue and test all the cathode channels. All of the bad channels (with few possible exceptions) are in the third and fourth package, where we didn't use the vacuum fixture for the cathodes. Eugene suggested to have a special session to discuss the lessons learned from the FDC production and installation.

- Beni and Cody made significant progress in reading the fADC125. After re-arranging the modules they were able to read half of the crates. Then Cody changed the VME clock from 80MHz to 40MHz and they were able to read the other half of the crates, as well. Still the results coming from the DAQ are with low statistics. They are enough to see that there are no significant problems (like bad connector) but most of the channels have few events and some have zeroes, so one can't count how many are the bad channels. Many of the channels on the most upstream cell (next to the mesh) have noise (up to 20kHz) that results in big output files. Beni will implement individual thresholds for all the channels to suppress the noisy channels.

- All the wire channels are working fine as tested with the DAQ. From time to time there's noise (up to few kHz) on some channels, again in the first package.


- The FDC was put in its final position based on marks on the rails. After the survey we may not need to move the FDC if it is within 2-3 mm from the nominal (Eugene). The most important is to install the stops before the survey to make sure FDC doesn't move and we know its position.

- Dave and Bill put a lot of efforts on Monday to make fixing tools. For the future, different ideas wre discussed. Bill: the simplest thing would be to have a spring that is compressed when inserted and then released. Bill could not find a telescopic mechanism with such sizes, but such tool can be made.

- Again we had long discussions about the cathode scan and possible systematic errors in the measurements. The reason for the jump of 250 microns in the middle is still unknown. People from Hexagon confirmed it is possible to have deviations after re-homing of the device, but not so big (Bill). We discussed if we want to scan another cathode. Eugene prefers to use different method, like using simple ruler, to check the scan results. Lubomir will do additional analyses to check the consistency of the measurements.


- FDC and the mesh have to be grounded before we can discuss the noise problems. We may need improvement of the ground of the most upstream cell, next to the mesh.

- A LV cable was cut during installation: the insulation and 3 out of 10 wires were damaged, but there's still LV on the cards. Chris ordered parts needed to fix it. Most likely it was cut on the rail shims; we were sliding the cables on the top of the rails without realizing how sharp the shims are! Fernando suggested inspecting all the cables for such cuts.

- Several 3M connectors have to be fixed, but we want to make sure they are really bad, because fixing is difficult. One connector can't be disconnected from the module: the plastic part separated from the metal base.

- 12 pre-amps that were replaced on the FDC were given to Nick for evaluation. Not all are necessarily bad, it could be also the connector on the pre-amp/daughter card that after replacement resulted in better connection.


- For the next three weeks Beni will be chairing the meetings; Lubomir is in vacation.